# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ 0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
Lemma 9 shows us that the construction of NE-simple predicates in [BREMER201215] can be used in the setting of relational patterns under abelian equivalence in exactly the same way. Mainly, this is caused by the fact that each variable $\hat{x}_{i,j}\in X_{s}$ is substituted by a unary string. By that, each related variable has to be substituted exactly in the same way as they are substituted in [BREMER201215], i.e. equally. Having shown that the same simple predicates from [BREMER201215] also work in this setting in exactly the same way, we define $\pi_{4}$ to $\pi_{\mu}$ exactly as they are defined in [BREMER201215], as all of those predicates are given described as simple predicates.
We set $\alpha_{1}:=\#\#enc_{NE}(I)\#\#x\#0^{6}0^{1}0\#\#$ and $\alpha_{2}:=y00$ for pairwise distinct variables $x$ and $y$ (notice, that $x$ and $y$ are the only variables in the pattern $\alpha$, hence for $(\alpha,r_{\alpha})$ we have $r_{\alpha}=\emptyset$). By the construction of $\pi_{1}$ to $\pi_{\mu}$ and essentially the same argument over the embedding of all invalid computations in $(\beta,r_{\beta})$ from [BREMER201215], we obtain that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$. This concludes the case $|\Sigma|=2$.
For the case of larger alphabets, i.e. alphabets of size $|\Sigma|\geq 3$, the exact construction in [BREMER201215] can also be adapted to the formality of our setting and works by the same arguments as in [BREMER201215]. The relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ need to be slightly adapted to work with larger alphabets. Essentially, suffixes to both the patterns $\alpha$ and $\beta$ are added to allow for another new variable $\hat{x}$ to be mapped to any letter of $\Sigma$ (similar to how we ensured that one variable $x_{i,j}$ is mapped to #\#). Then, $2$ new predicates are added to ensure that if this letter occurs in any substitution of $h(\alpha_{1})$ or $h(\alpha_{2})$, then one of these predicates is satisfied.
@@
where $y_{\mu+1,1},y_{\mu+1,2},y_{\mu+2},\hat{y}_{\mu+1},\hat{y}_{\mu+2,1},$ and $\hat{y}_{\mu+2,2}$ are new and pairwise distinct and unrelated variables. Remember, that we have $(r_{ab},\tilde{x}_{1,1},\tilde{x}_{1,2})\in r_{\tilde{\beta}}$ and $(r_{ab},\tilde{x}_{2,1},\tilde{x}_{2,2})\in r_{\tilde{\beta}}$. Now, using these constructions, we obtain the exact argument as in [BREMER201215] to obtain the final result. The main argument why this construction can be used by the same argument is that it works over unary substitutions of $\tilde{x}_{1,1},\tilde{x}_{1,2},\tilde{x}_{2,1},$ and $\tilde{x}_{2,2}$ of length $1$. Hence, using substitutions of this type, abelian equivalence behaves exactly the same as equality between variables. Thus, we omit the details at this point and refer to [BREMER201215] for a more detailed proof.
@@
0.B.5 Undecidability of the Inclusion Problem for Erasing Relational Pattern Languages under the Reversal Relation
Proof of Theorem˜4.1:
Proof
The proof is based on the construction by Freydenberger and Reidenbach [Freydenberger2010]. For better understanding for those who don’t know the construction and for a really detailed verification, we copied the proof of Theorem 2 of [Freydenberger2010] and adapted it at several points. The main difference is the definition of $u$ and the definition of the delimiter. Why these adaptations work is shorter described in the proof of Theorem˜4.2 for those who already are familiar with the construction. The adaptations of Theorem˜4.1 and Theorem˜4.2 are very similar.
@@
with, for all $i\in\{1,\ldots,\mu\}$, $\hat{\beta}_{i}:=x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}}$ and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{1}},x_{1_{2}},x_{1_{3}},x_{1_{4}},x_{1_{5}},\ldots,x_{\mu},x_{\mu^{\prime}},x_{\mu_{1}},x_{\mu_{2}},x_{\mu_{3}},x_{\mu_{4}},x_{\mu_{5}}$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all
$\gamma_{i},\delta_{i},\eta_{i}\in X^{*}$ are terminal-free patterns. The patterns $\gamma_{i}$ and $\delta_{i}$ shall be defined later; for now, we only mention:
@@
Thus, for every $i$, the elements of var($\gamma_{i}\delta_{i}\eta_{i}$) appear nowhere but in these three factors. Let $H$ be the set of all substitutions $\sigma:(\Sigma\ \cup\ \{x,y\})^{*}\rightarrow\Sigma^{*}$. We interpret each triple ($\gamma_{i},\delta_{i},\eta_{i}$) as a predicate $\pi_{i}:H\rightarrow\{0,1\}$ in such a way that $\sigma\in H$ satisfies $\pi_{i}$ if there exists a morphism $\tau:\text{var}(\gamma_{i}\delta_{i}\eta_{i})^{*}\rightarrow\Sigma^{*}$ with $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ - in the terminology of word equations (cf. Karhumäki et al. [DBLP:journals/jacm/KarhumakiMP00]), this means that $\sigma$ satisfies $\pi_{i}$ if and only if the system consisting of the three equations $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ has a solution $\tau$. Later, we shall see that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})$ exactly contains those $\sigma(\alpha_{A})$ for which $\sigma$ does not satisfy any of $\pi_{1}$ to $\pi_{\mu}$, and choose these predicates to describe $\mathtt{InvalC}(A)$. The encoding of $\mathtt{InvalC}(A)$ shall be handled by $\pi_{4}$ to $\pi_{\mu}$, as each of these predicates describes a sufficient criterium for membership in $\mathtt{InvalC}(A)$. But at first we need a considerable amount of technical preparations. A substitution $\sigma$ is of good form if $\sigma(x)\in\{0,\#\}^{*}$, $\sigma(x)$ does not contain $^{3}$ as a factor, and $\sigma(y)\in 0^{*}$. Otherwise, $\sigma$ is of bad form. The predicates $\pi_{1}$ and $\pi_{2}$ handle all cases where $\sigma$ is of bad form and are defined through
@@
where $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2},\hat{z}_{1_{6}},\hat{z}_{1_{7}},\hat{z}_{1_{8}}$ and $\hat{z}_{2_{6}}$ are pairwise distinct variables with $(\hat{z}_{1_{1}},\hat{z}_{1_{6}}),(\hat{z}_{1_{1}},\hat{z}_{1_{7}}),(\hat{z}_{1_{1}},\hat{z}_{1_{8}}),(\hat{z}_{2_{1}},\hat{z}_{2_{6}})\in R$. Recall that
$\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ for all $i$. It is not very difficult to see that $\pi_{1}$ and $\pi_{2}$ characterize the morphisms that are of bad form:
Lemma 10
@@
Analogously, if $\sigma$ satisfies $\pi_{2}$, then $\sigma(y)$ contains the letter #\#, and $\sigma$ is of bad form.
@@
As $\sigma$ is of good form, $\sigma(0\#^{3}0\ x\ 0\#^{3}0\ y\ 0\#^{3}0)$ contains exactly three occurrences of $^{3}$. But there are already three occurrences of $^{3}$ in $\tau(\hat{\beta}_{i})=0\#^{3}0\ \tau(\gamma_{i})\ 0\#^{3}0\ \tau(\delta_{i})\ 0\#^{3}0$. This, and Equation˜5, lead to $\tau(\hat{\beta}_{j})=\epsilon$ for all $j\neq i$ and, more importantly, $\tau(\gamma_{i})=\sigma(x)$ and $\tau(\delta_{i})=\sigma(y)$. Therefore, $\sigma$ satisfies the predicate $\pi_{i}$.
@@
where all of $\gamma_{3,1}$ to $\gamma_{3,4}$ and $\hat{\gamma}_{3,1}$ to $\hat{\gamma}_{3,3}$ and $\hat{\gamma}_{3,1^{\prime}}$ to $\hat{\gamma}_{3,3^{\prime}}$ are pairwise distinct variables with $(\hat{\gamma}_{3,1},\hat{\gamma}_{3,1^{\prime}}),(\hat{\gamma}_{3,2},\hat{\gamma}_{3,2^{\prime}}),(\hat{\gamma}_{3,3},\hat{\gamma}_{3,3^{\prime}})\in R$. Clearly, if some $\sigma\in H$ satisfies $\pi_{3}$, $\sigma(y)$ is a concatenation of three (possibly empty) reversed factors of $\sigma(x)$. Thus, if $\sigma$ satisfies none of $\pi_{1}$ to $\pi_{3}$, $\sigma(y)$ must be longer than the three longest non-overlapping sequences of 0s in $\sigma(x)$. This allows us to identify a class of predicates definable by a rather simple kind of expression, which we use to define $\pi_{4}$ to $\pi_{\mu}$ in a less technical way. The following definitions are adapted to work with our setting of relational patterns.
@@
Proof
We first consider the case of $L_{1}=L_{2}=\Sigma^{*}$. Assume that $\pi_{S}$ is a simple predicate, and $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ is a pattern such that $\sigma\in G_{v_{R}}$ satisfies $\pi_{S}$ if and only if $\sigma(x)\in L_{1}\ S(p,v_{R})\ L_{2}$. Then define $\gamma:=y_{1}p^{\prime}y_{2}$, where $p^{\prime}$ is obtained from $p$ by replacing all $k,k\in\mathbb{N}_{0}$ occurrences of 0 with new variables $\tilde{z}_{j},j\in[k]$ with $(\tilde{z}_{1},\tilde{z}_{j})\in R$ for all $j\in[k]$ and all $g,g\in\mathbb{N}_{0}$ occurrences of #\# with different variables $\ddot{z}_{j},j\in[g]$ with $(\ddot{z}_{1},\ddot{z}_{j})\in R$ for all $j\in[g]$, while leaving all present elements of $X^{\prime}$ unchanged. Furthermore, let $\delta:=\hat{x}_{1,j}\hat{x}_{2,j}\hat{x}_{3,j}\hat{y}$ with $j\in\mathbb{N}$ such that $\hat{x}_{1,j},\hat{x}_{2,j},\hat{x}_{3,j}$ are new variables and (in order to stay consistent with the $\eta_{i}$ appearing in $\beta_{A}$) $\eta:=\tilde{z}\ \ddot{z}_{g+1}\ \ddot{z}_{g+2}\ \ddot{z}_{g+3}\ \ddot{z}_{g+4}\ \ddot{z}_{g+5}\ \tilde{z}^{\prime}$ with $(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$.
@@
$(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$ and u=0#####0u=0\#\#\#\#\#0, $\tau(\tilde{z})=\tau(\tilde{z}^{\prime})=0$ and $\tau(\ddot{z}_{j})=\#$ for all $j\in\{g+1,\ldots,g+5\}$ must hold. By definition $\tau(y_{1}),\tau(y_{2})\in\Sigma^{*}$. If we define $\rho(\hat{x}_{i,j}):=\tau(\hat{x}_{i,j})$ for all $i\in\{1,2,3\},j\in\mathbb{N}$ such that $\tau(\hat{x}_{i,j})$ is defined and $\rho(\hat{x}_{i,j}):=\rho(\hat{x}_{i,j^{\prime}})$ for all other $i\in\{1,2,3\},j\in\mathbb{N}$ for one arbitrary, specific $j^{\prime}\in\mathbb{N}$ such that $\rho(\hat{x}_{i,j^{\prime}})$ is defined, we see that $\sigma(x)\in L_{1}S(p,v_{R})L_{2}$ holds. Thus, $\sigma$ satisfies $\pi_{S}$ as well.
The other three cases for choices of $L_{1}$ and $L_{2}$ can be handled analogously by omitting $y_{1}$ or $y_{2}$ as needed. Note that this proof also works in the case $p=\epsilon$.
@@
Regarding the only if direction, assume some $w\in\Sigma^{*}$ is of good structure. Define $\sigma$ by $\sigma(x)=w$ and $\sigma(y)=0^{|w|+1}$. As $\sigma$ is of good form, Lemma˜10 demonstrates that $\sigma$ satisfies neither $\pi_{1}$ nor $\pi_{2}$; and as $\sigma(y)$ is longer than any word which results from concatenating any number of non-overlapping factors of the form $0^{*}$ of $w$, $\pi_{3}$ cannot be satisfied either. By looking at the cases used above to define $\pi_{4}$ to $\pi_{13}$, we see that none of these predicates can be satisfied.
@@
We conclude that $A$ has an accepting computation iff $L_{E}(\alpha_{A},r_{\alpha_{A}})$ is not a subset of $L_{E}(\beta_{A},r_{\beta_{A}})$. Therefore, any algorithm deciding the inclusion problem of Theorem˜4.1 can be used to decide whether a nondeterministic 2-counter automaton without input has an accepting computation. As this problem is known to be undecidable, the inclusion problem of Theorem˜4.1 is also undecidable.
@@
$(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}}),(\ddot{z}_{i,k_{1}},\ddot{z}_{i,k_{2}})\in R.$

### v2

@@ 0 $(0,R,q_{1})$ $(1,L,q_{11})$ $(0,R,q_{12})$ $(0,R,q_{13})$ $(0,L,q_{2})$ $(0,L,q_{3})$ $(0,R,q_{14})$
Lemma 9 shows us that the construction of NE-simple predicates in [5] can be used in the setting of relational patterns under abelian equivalence in exactly the same way. Mainly, this is caused by the fact that each variable $\hat{x}_{i,j}\in X_{s}$ is substituted by a unary string. By that, each related variable has to be substituted exactly in the same way as they are substituted in [5], i.e. equally. Having shown that the same simple predicates from [5] also work in this setting in exactly the same way, we define $\pi_{4}$ to $\pi_{\mu}$ exactly as they are defined in [5], as all of those predicates are given described as simple predicates.
We set $\alpha_{1}:=\#\#enc_{NE}(I)\#\#x\#0^{6}0^{1}0\#\#$ and $\alpha_{2}:=y00$ for pairwise distinct variables $x$ and $y$ (notice, that $x$ and $y$ are the only variables in the pattern $\alpha$, hence for $(\alpha,r_{\alpha})$ we have $r_{\alpha}=\emptyset$). By the construction of $\pi_{1}$ to $\pi_{\mu}$ and essentially the same argument over the embedding of all invalid computations in $(\beta,r_{\beta})$ from [5], we obtain that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$. This concludes the case $|\Sigma|=2$.
For the case of larger alphabets, i.e. alphabets of size $|\Sigma|\geq 3$, the exact construction in [5] can also be adapted to the formality of our setting and works by the same arguments as in [5]. The relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ need to be slightly adapted to work with larger alphabets. Essentially, suffixes to both the patterns $\alpha$ and $\beta$ are added to allow for another new variable $\hat{x}$ to be mapped to any letter of $\Sigma$ (similar to how we ensured that one variable $x_{i,j}$ is mapped to #\#). Then, $2$ new predicates are added to ensure that if this letter occurs in any substitution of $h(\alpha_{1})$ or $h(\alpha_{2})$, then one of these predicates is satisfied.
@@
where $y_{\mu+1,1},y_{\mu+1,2},y_{\mu+2},\hat{y}_{\mu+1},\hat{y}_{\mu+2,1},$ and $\hat{y}_{\mu+2,2}$ are new and pairwise distinct and unrelated variables. Remember, that we have $(r_{ab},\tilde{x}_{1,1},\tilde{x}_{1,2})\in r_{\tilde{\beta}}$ and $(r_{ab},\tilde{x}_{2,1},\tilde{x}_{2,2})\in r_{\tilde{\beta}}$. Now, using these constructions, we obtain the exact argument as in [5] to obtain the final result. The main argument why this construction can be used by the same argument is that it works over unary substitutions of $\tilde{x}_{1,1},\tilde{x}_{1,2},\tilde{x}_{2,1},$ and $\tilde{x}_{2,2}$ of length $1$. Hence, using substitutions of this type, abelian equivalence behaves exactly the same as equality between variables. Thus, we omit the details at this point and refer to [5] for a more detailed proof.
@@
0.B.2 Proof of Lemma˜4
Proof.
The proof relies on the proof of Lemma˜3.
(1:) As shown in the first part of proof of Lemma˜3, the language of accepted words of $\alpha$ is
$L_{E}(\alpha,r_{\alpha})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}$
and the language of accepted words of $\beta$ is
$L_{E}(\beta,r_{\beta})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
Thus, $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$ if and only if $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$. This property can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
(2:) Assume $|\Sigma|=1$ and $a\in\Sigma$. In order to be included, $\alpha$ must have at least as many terminals as $\beta$. Otherwise we could set all variables in $\alpha$ to $\varepsilon$ and obtain a word that is not in the language of $\beta$. Let $\ell_{\alpha}$ be the number of terminals in $\alpha$ and $\ell_{\beta}$ be the number of terminals in $\beta$. As shown in the second part of proof of Lemma˜3, the language of accepted words of $\alpha$ is
$L_{E}(\alpha,r_{\alpha})=\{a^{\ell_{\alpha}}\}\cup\{w\mid|w|\geq\ell_{\alpha}+\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}$
and the language of accepted words of $\beta$ is
$L_{E}(\beta,r_{\beta})=\{a^{\ell_{\beta}}\}\cup\{w\mid|w|\geq\ell_{\beta}+\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
Thus, $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$ if and only $\ell_{\alpha}\geq\ell_{\beta}$ and $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$. This can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes. ∎
0.B.3 Proof Sketch of Theorem˜4.3
0.B.3.1 (Nondeterministic 2-Counter Automata Without Input):
A nondeterministic 2-counter automaton without input (see e.g. [21]) is a 4-tuple $A=(Q,\delta,q_{0},F)$ which consists of a set of states $Q$, a transition function $\delta:Q\times\{0,1\}^{2}\rightarrow\mathcal{P}(Q\times\{\-1,0,+1\}^{2})$, an initial state $q_{0}\in Q$, and a set of accepting states $F\subseteq Q$. A configuration of $A$ is defined as a triple $(q,m_{1},m_{2})\in Q\times\mathbb{N}\times\mathbb{N}$ in which $q$ indicates the current state and $m_{1}$ and $m_{2}$ indicate the contents of the first and second counter. We define the relation $\vdash_{A}$ on $Q\times\mathbb{N}\times\mathbb{N}$ by $\delta$ as follows. For two configurations $(p,m_{1},m_{2})$ and $(q,n_{1},n_{2})$ we say that $(p,m_{1},m_{2})\vdash_{A}(q,n_{1},n_{2})$ if and only if there exist $c_{1},c_{2}\in\{0,1\}$ and $r_{1},r_{2}\in\{-1,0,+1\}$ such that
  1. 1.
    if $m_{i}=0$ then $c_{i}=0$, otherwise if $m_{i}>0$, then $c_{i}=1$, for $i\in\{1,2\}$,
  2. 2.
    $n_{i}=m_{i}+r_{i}$ for $i\in\{1,2\}$,
  3. 3.
    $(q,r_{1},r_{2})\in\delta(p,c_{1},c_{2})$, and
  4. 4.
    we assume if $c_{i}=0$ then $r_{i}\neq-1$ for $i\in\{1,2\}$.
and let $\mathtt{InvalC}(A)=\{0,\#\}^{*}\setminus\mathtt{ValC}(A)$. The emptiness problem for deterministic 2-counter-automata is undecidable (cf. e.g. [21, 29]), thus it is also undecidable whether a nondeterministic 2-counter automaton without input has an accepting computation [13, 23]. That the emptiness problem for universal Turing machines is undecidable is a known fact.
0.B.3.2 (The Full Proof Idea for $r_{rev}$):
Proof.
The proof is based on the construction by Freydenberger and Reidenbach [13]. For better understanding for those who don’t know the construction and for a really detailed verification, we copied the proof of Theorem 2 of [13] and adapted it at several points. The main difference is the definition of $u$ and the definition of the delimiter. Why these adaptations work is shorter described in the proof part for $r_{ab}$ for those who already are familiar with the construction. The adaptations for $r_{rev}$ and $r_{ab}$ are very similar. For now, assume $R=\{r_{rev}\}$.
@@ 0.B.2 Proof of Lemma˜4
with, for all $i\in\{1,\ldots,\mu\}$, $\hat{\beta}_{i}:=x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}}$ and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{1}},x_{1_{2}},x_{1_{3}},x_{1_{4}},x_{1_{5}},\ldots,x_{\mu},x_{\mu^{\prime}},x_{\mu_{1}},x_{\mu_{2}},x_{\mu_{3}},x_{\mu_{4}},x_{\mu_{5}}$ are distinct variables with $(r_{rev},x_{i},x_{i^{\prime}})$, $(r_{rev},x_{i},x_{i_{1}})$, $(r_{rev},x_{i},x_{i_{2}})$, $(r_{rev},x_{i},x_{i_{3}})$, $(r_{rev},x_{i},x_{i_{4}})$, $(r_{rev},x_{i},x_{i_{5}})\in r_{\beta_{A}}$ and all $\gamma_{i},\delta_{i},\eta_{i}\in X^{*}$ are terminal-free patterns. The patterns $\gamma_{i}$ and $\delta_{i}$ shall be defined later; for now, we only mention:
@@
Thus, for every $i$, the elements of var($\gamma_{i}\delta_{i}\eta_{i}$) appear nowhere but in these three factors. Let $H$ be the set of all substitutions $\sigma:(\Sigma\ \cup\ \{x,y\})^{*}\rightarrow\Sigma^{*}$. We interpret each triple ($\gamma_{i},\delta_{i},\eta_{i}$) as a predicate $\pi_{i}:H\rightarrow\{0,1\}$ in such a way that $\sigma\in H$ satisfies $\pi_{i}$ if there exists a morphism $\tau:\text{var}(\gamma_{i}\delta_{i}\eta_{i})^{*}\rightarrow\Sigma^{*}$ with $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ - in the terminology of word equations (cf. Karhumäki et al. [24]), this means that $\sigma$ satisfies $\pi_{i}$ if and only if the system consisting of the three equations $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ has a solution $\tau$. Later, we shall see that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})$ exactly contains those $\sigma(\alpha_{A})$ for which $\sigma$ does not satisfy any of $\pi_{1}$ to $\pi_{\mu}$, and choose these predicates to describe $\mathtt{InvalC}(A)$. The encoding of $\mathtt{InvalC}(A)$ shall be handled by $\pi_{4}$ to $\pi_{\mu}$, as each of these predicates describes a sufficient criterium for membership in $\mathtt{InvalC}(A)$. But at first we need a considerable amount of technical preparations. A substitution $\sigma$ is of good form if $\sigma(x)\in\{0,\#\}^{*}$, $\sigma(x)$ does not contain $^{3}$ as a factor, and $\sigma(y)\in 0^{*}$. Otherwise, $\sigma$ is of bad form. The predicates $\pi_{1}$ and $\pi_{2}$ handle all cases where $\sigma$ is of bad form and are defined through
@@
where $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2},\hat{z}_{1_{6}},\hat{z}_{1_{7}},\hat{z}_{1_{8}}$ and $\hat{z}_{2_{6}}$ are pairwise distinct variables with $(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{6}}),(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{7}}),(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{8}}),(r_{rev},\hat{z}_{2_{1}},\hat{z}_{2_{6}})\in r_{\beta_{A}}$. Recall that $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ for all $i$. It is not very difficult to see that $\pi_{1}$ and $\pi_{2}$ characterize the morphisms that are of bad form:
Lemma 10.
@@
Analogously, if $\sigma$ satisfies $\pi_{2}$, then $\sigma(y)$ contains the letter #\#, and $\sigma$ is of bad form. ∎
@@
As $\sigma$ is of good form, $\sigma(0\#^{3}0\ x\ 0\#^{3}0\ y\ 0\#^{3}0)$ contains exactly three occurrences of $^{3}$. But there are already three occurrences of $^{3}$ in $\tau(\hat{\beta}_{i})=0\#^{3}0\ \tau(\gamma_{i})\ 0\#^{3}0\ \tau(\delta_{i})\ 0\#^{3}0$. This, and Equation˜5, lead to $\tau(\hat{\beta}_{j})=\epsilon$ for all $j\neq i$ and, more importantly, $\tau(\gamma_{i})=\sigma(x)$ and $\tau(\delta_{i})=\sigma(y)$. Therefore, $\sigma$ satisfies the predicate $\pi_{i}$. ∎
@@
where all of $\gamma_{3,1}$ to $\gamma_{3,4}$ and $\hat{\gamma}_{3,1}$ to $\hat{\gamma}_{3,3}$ and $\hat{\gamma}_{3,1^{\prime}}$ to $\hat{\gamma}_{3,3^{\prime}}$ are pairwise distinct variables with $(r_{rev},\hat{\gamma}_{3,1},\hat{\gamma}_{3,1^{\prime}}),(r_{rev},\hat{\gamma}_{3,2},\hat{\gamma}_{3,2^{\prime}}),(r_{rev},\hat{\gamma}_{3,3},\hat{\gamma}_{3,3^{\prime}})\in r_{\beta_{A}}$. Clearly, if some $\sigma\in H$ satisfies $\pi_{3}$, $\sigma(y)$ is a concatenation of three (possibly empty) reversed factors of $\sigma(x)$. Thus, if $\sigma$ satisfies none of $\pi_{1}$ to $\pi_{3}$, $\sigma(y)$ must be longer than the three longest non-overlapping sequences of 0s in $\sigma(x)$. This allows us to identify a class of predicates definable by a rather simple kind of expression, which we use to define $\pi_{4}$ to $\pi_{\mu}$ in a less technical way. The following definitions are adapted to work with our setting of relational patterns.
@@
Proof.
[…]
@@
The other three cases for choices of $L_{1}$ and $L_{2}$ can be handled analogously by omitting $y_{1}$ or $y_{2}$ as needed. Note that this proof also works in the case $p=\epsilon$. ∎
@@
Regarding the only if direction, assume some $w\in\Sigma^{*}$ is of good structure. Define $\sigma$ by $\sigma(x)=w$ and $\sigma(y)=0^{|w|+1}$. As $\sigma$ is of good form, Lemma˜10 demonstrates that $\sigma$ satisfies neither $\pi_{1}$ nor $\pi_{2}$; and as $\sigma(y)$ is longer than any word which results from concatenating any number of non-overlapping factors of the form $0^{*}$ of $w$, $\pi_{3}$ cannot be satisfied either. By looking at the cases used above to define $\pi_{4}$ to $\pi_{13}$, we see that none of these predicates can be satisfied. ∎
@@
We conclude that $A$ has an accepting computation iff $L_{E}(\alpha_{A},r_{\alpha_{A}})$ is not a subset of $L_{E}(\beta_{A},r_{\beta_{A}})$. Therefore, any algorithm deciding the inclusion problem under $r_{rev}$ can be used to decide whether a nondeterministic 2-counter automaton without input has an accepting computation. As this problem is known to be undecidable, the inclusion problem under $r_{rev}$ is also undecidable.
@@
$(r_{rev},z_{i},z_{i^{\prime}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{3}}),$
$(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{5}}),(r_{rev},\ddot{z}_{i,k_{1}},\ddot{z}_{i,k_{2}})\in r_{\beta_{A}}.$